# Bright traveling breathers in media with long-range, nonconvex dispersion

### v2

@@ Appendix A Validation of the NCG algorithm on mKdV breathers
$u_{t}+3u^{2}u_{x}+u_{xxx}=0,$ (33)
is known to possess bright breathers with closed form expressions. We assess the performance of the NCG algorithm on recovering the mKdV breathers situated on a zero background. The expression for the two-parameter family of zero-mean mKdV breathers in the envelope reference frame $u(x,t)=\beta(x-ct,t)$ is Wadati (1973).
@@
Figure 12: The continuation path (black dots) marked out in the $c$-$\Omega$ phase plane; in the weakly nonlinear regime the path approaches the linear dispersion curve (the magenta curve: $\Omega_{0}(c_{0})$) as $c\to-3$. The inset shows slices at $\tau=0$ of the computed (a) weakly nonlinear mKdV breather and a (b) strongly nonlinear breather.
There are two distinguished limits of breather solutions to the mKdV equation (a) the weakly nonlinear, NLS regime (${\kappa_{1}}/{\kappa_{2}}\ll 1$) and (b) the strongly nonlinear regime (${\kappa_{1}}/{\kappa_{2}}\sim\mathcal{O}(1)$) where breathers tend to double-pole solutions Wadati and Ohkuma (1982). For $T=(2\pi/\Omega)=3.1105$, where $\Omega$ is the angular frequency of oscillations in the envelope reference frame, we recover both the distinguished limits using the Newton conjugate gradient algorithm, coupled to a $c$-continuation line search. Our computational parameters are: $\Delta x=0.05$, $\Delta t\approx 0.1$ and the computational spatial domain $4002L=$ 400. Along the entire constant $T$ path on the $c-\Omega$ plane, the space-time infinity norm of the error in the numerical solution is kept approximately $\sim 10^{-7}$. The results of numerical continuation are summarized in Fig. 12.

### v3

@@ Appendix A Validation of the NCG algorithm on mKdV breathers
$u_{t}+3u^{2}u_{x}+u_{xxx}=0$ (33)
is known to possess bright breathers with closed-form expressions. We assess the performance of the NCG algorithm on recovering the mKdV breathers situated on a zero background. The expression for the two-parameter family of zero-mean mKdV breathers in the envelope reference frame $u(x,t)=\beta(x-ct,t)$ is Wadati (1973)
@@
Figure 12: Continuation path (black dots) marked in the $c$-$\Omega$ phase plane; in the weakly nonlinear regime the path approaches the linear dispersion curve [magenta curve: $\Omega_{0}(c_{0})$] as $c\to-3$. The inset shows slices at $\tau=0$ of (a) the computed weakly nonlinear mKdV breather and (b) a strongly nonlinear breather.
There are two distinguished limits of breather solutions to the mKdV equation (a) the weakly nonlinear NLS regime (${\kappa_{1}}/{\kappa_{2}}\ll 1$) and (b) the strongly nonlinear regime [${\kappa_{1}}/{\kappa_{2}}\sim{O}(1)$] where breathers tend to double-pole solutions Wadati and Ohkuma (1982). For $T=2\pi/\Omega=3.1105$, where $\Omega$ is the angular frequency of oscillations in the envelope reference frame, we recover both the distinguished limits using the Newton conjugate gradient algorithm, coupled to a $c$-continuation line search. Our computational parameters are $\Delta x=0.05$, $\Delta t\approx 0.1$, and the computational spatial domain $4002L=$ 400. Along the entire constant $T$ path on the $c$-$\Omega$ plane, the space-time infinity norm of the error in the numerical solution is kept approximately $10^{-7}$. The results of numerical continuation are summarized in Fig. 12.